OpenStudy (anonymous):

Solve the following system of equations: 3x - 2y = 6 6x - 4y = 14 (0, 0) (6, 14) Infinitely many solutions No solutions

OpenStudy (anonymous):

Okay, are you using substitution or elimination ?

OpenStudy (anonymous):

Substitution I think

OpenStudy (anonymous):

To solve the system, you must eliminate either the x or the y & then solve. I choose the x. To get the x's to cancel out, multiply the top equation by -2 -2(3x-2y=6) -6x+4y=-12 6x-4y=14 ____________ 0x+0y=-2 \[0\neq-2\]

OpenStudy (anonymous):

so...no solution.

OpenStudy (anonymous):

Haha thx DSS:)